OpenStudy (anonymous):

Given the function f(x) = 1/4x, find the value of f(2). Im not sure if my answers are correct.

OpenStudy (lgbasallote):

just sub in 2 f(2) = 14(2) is that what you go?

OpenStudy (anonymous):

a.) 1/4 b.) 2 c.) 1 d.) 1.2

OpenStudy (lgbasallote):

got*

OpenStudy (lgbasallote):

1/4(2)

OpenStudy (anonymous):

so thats the one thats correct... hmmm.

OpenStudy (lgbasallote):

wait is that \[\frac{1}{4x}\] ot \[\frac{1}{4}x\]

OpenStudy (lgbasallote):

or*

OpenStudy (anonymous):

@lgbasallote i got the same 1/4(2) which is 1/8

OpenStudy (anonymous):

its just 1/4

OpenStudy (anonymous):

Oh okay are you still confused if its 4x?

OpenStudy (anonymous):

1/4x**

OpenStudy (lgbasallote):

oh then it becomes \[f(2) = \frac{1}{4} (2) = \frac{2}{4}\] can you simplify that?

OpenStudy (anonymous):

into 1/2??

OpenStudy (lgbasallote):

yep..is that in the choices?

OpenStudy (anonymous):

1.2

OpenStudy (anonymous):

aha

OpenStudy (lgbasallote):

is 1.2 supposed to be 1/2?

OpenStudy (anonymous):

i doubt it.

OpenStudy (anonymous):

@srosenfeld0938 you need to start writing your questions better.. x should be writen the write way.. the answer is 2/4 which is?

OpenStudy (lgbasallote):

wait...what is it again? is it \[\frac{1}{4x}\] or \[\frac{1}{4} x\] just say first or second

OpenStudy (anonymous):

1/2

OpenStudy (lgbasallote):

what? im asking what the question is

OpenStudy (anonymous):

first

OpenStudy (lgbasallote):

first...

OpenStudy (lgbasallote):

then \[f(2) = \frac{1}{4(2)} = \frac{1}{8}\]that should be the right one o.O